# Learning-based Power Control for Secure Covert Semantic Communication

### v1

@@ V Conclusion
In this paper, we proposed a learning-based approach to optimal power control for the proposed covert SemCom, aimed at resolving the security dilemmas inherent in SemCom. Specifically, we introduced a power regulator and a friendly jammer to achieve covert communication within the context of SemCom. To address the increased energy usage introduced by covert SemCom, we formulated power control as an optimization problem and utilized the SAC algorithm to optimize power control between the transmitter and the friendly jammer, which strikes a balance between the effectiveness of covert transmission and SemCom. Numerical results have demonstrated the effectiveness of the proposed optimization strategy within the covert SemCom framework. For future work, we will explore ways to integrate the concept of the Mixture of Experts (MoE) system into our covert SemCom framework, ensuring reliable communication across varied network conditions.

### v2

@@ V Conclusion
In this paper, we have proposed a learning-based approach to optimal power control for the proposed covert SemCom, aimed at resolving the security dilemmas inherent in SemCom. Specifically, we have introduced a power regulator and a friendly jammer to achieve covert communication within the context of SemCom. To address the increased energy usage introduced by covert SemCom, we have formulated power control as an optimization problem and utilized the SAC algorithm to optimize power control between the transmitter and the friendly jammer, which strikes a balance between the effectiveness of covert transmission and SemCom. Numerical results have demonstrated the effectiveness of the proposed optimization strategy within the covert SemCom framework. For future work, we will explore ways to integrate the concept of the Mixture of Experts (MoE) system into our covert SemCom framework, ensuring reliable communication across varied network conditions.